OpenStudy (anonymous):

A cruise ship has circular windows in its state rooms. What is the circumference of one of these windows if it's diameter is 15 inches?

OpenStudy (anonymous):

circumference = \(2\pi r\)

OpenStudy (anonymous):

r = radius diameter = 2 \(\times\) radius

OpenStudy (anonymous):

how do u find radius in this problem @gypsy1274

OpenStudy (anonymous):

They gave you the diameter, I have you the relationship between the diameter and the radius. solve for radius.

OpenStudy (anonymous):

do you do 15/2=7.5=radius?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so 7.5 is the radius so 2pi7.5^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

circumference = \(\pi \times 2 \times r\)

OpenStudy (anonymous):

so =353.43

jhonyy9 (jhonyy9):

not r squared just simple r

OpenStudy (anonymous):

oh ohk thanks

jhonyy9 (jhonyy9):

yw